Please factorize the polynomial (x+y)^2+3xy^2-x^3+xy-y^3+3x^2y
1
Expert's answer
2012-06-19T09:19:31-0400
If we take g1=x+y and g2=xy then our expression can be simplified as
(g1)^3+3(g1)(g2)+g2-(g1)^3+3(g1)(g2)=(6g1+1)g2=
=(6x+6y+1)xy